OpenStudy (anonymous):

A certain radioactive material is known to decay at a rate proportional to the amount present. If after one hour, it is observed that 10% of the material has decayed, find the half-life of the material.

OpenStudy (anonymous):

Can someone help me solve this?

OpenStudy (anonymous):

do you have answer choices??

OpenStudy (anonymous):

Let initial amount of radioactive material is No then N(0)=No , N(1)=(1−0.1)No=0.9No .

OpenStudy (anonymous):

i have answer. It is 6.6hrs

OpenStudy (anonymous):

We need to find such t that N(t)=No/2 .

OpenStudy (anonymous):

you are correct ^-^ @BizPro

OpenStudy (anonymous):

hi again @BizPro

OpenStudy (anonymous):

no sarah, that's the answer given, i'm just not sure how to do it.

OpenStudy (anonymous):

ok Im tellin you step by step . got first two step ???

OpenStudy (anonymous):

why No?

OpenStudy (anonymous):

or is that just short for number?

OpenStudy (anonymous):

we are supposing Let initial amount of radioactive material is No then N(0)=No , N(1)=(1−0.1)No=0.9No . We need to find such t that N(t)=No/2 . So, as was shown above, if N(0)=No then N(t)=Noe^kt . Since N(1)=0.9No then 0.9No=Noe^k⋅1 or k=ln(0.9) . Now equation has form N(t)=Noe^ln(0.9)t . Now, find t, such that N(t)=0.5No : 0.5No=Noe^ln(0.9)t or t=ln(0.5)/ln(0.9)≈6.58 hours.

OpenStudy (anonymous):

im sorry just a bit confused. Need read through it again.

OpenStudy (anonymous):

you know na that half-life is period of time it takes for the amount of material to decrease by half . So 10% or .1 is halfed then 5% or .5

OpenStudy (anonymous):

Oh Noe^kt is like Ae^kt

OpenStudy (anonymous):

No is initial amount that we suppose .

OpenStudy (anonymous):

These questions always confuse me :(

OpenStudy (anonymous):

initial amount is 1 , how much decayed as mentioned in question 10 % means 10/100 = .1 So subtract .1 from 1(initial amount) we get .9 left

OpenStudy (anonymous):

yeah i understand up till Noe^kt

OpenStudy (anonymous):

then u need to know this how to find growth and decay formula . Let N(t) denote amount of substance (or population) that is either growing or decaying. If we assume that dN/dt , the time rate of change of this amount of substance, is proportional to the amount of substance present, then dN/dt=kN where k is the constant of proportionality. When k is positive, population grows, when k is negative - population decays. We are assuming that N(t) is a differentiable, hence continuous, function of time. For population problems, where N(t) is actually discrete and integer-valued, this assumption is incorrect. Nonetheless, above model still provides a good approximation to she physical laws governing such a system. This differential equation is separable. Rewriting it we obtain dN/N=kdt . Integration of both sides gives: ln(N)=kt+c1 or N=Ce^kt where C=e^c1 . If we are given that N(0)=No thenNo=Ce^k⋅0 or C=no . So, N=Noe^kt .

OpenStudy (anonymous):

If still you dun get it then .-.

OpenStudy (anonymous):

Oh ok thnx :)

OpenStudy (anonymous):

yw ^-^